A 1.15-kΩkΩ resistor and a 525-mHmH inductor are connected in series to a 1400-HzHz generator with an rms voltage of 12.1 VV .

A) What is the rms current in the circuit?

B) What capacitance must be inserted in series with the resistor and inductor to reduce the rms current to half the value found in part A?
Answer #1

4] W- 27711400) - 28001 XL = Lw 3 R = 1.15X10 sh. 1150 2. *1= 525x16” x 28000 = 4618-14 I v z=1 R²+x2 = 4759.17 I I = V 1201